Could someone answer this problem : b+m=8 and 51b+8m=193. Please answer my queston.
Thank You,
Lincoln

Answer by Nate(3500)

I will tell you how to solve for b and m:
we will use substitution!
b+m=8 is the same thing as b=8-m
51b+8m=193--original equation
51(8-m)+8m=193--plug in 8-m for b
408-51m+8m=193--distribute
-43m=-215--get variable by itself
m=5--divide the equation by -43
use b+m=8 to find b
b+m=8--original equation
b+5=8--replace m with 5
b=3--subtract 5 from each side
b=3 and m=5 is your answer!